Suppose the total product (Q) is giving by the production functions; [q=4KL-0.5]
Where K= capital =12 and L= unit of labour -13.
Calculate the average product of labour
1
Expert's answer
2020-01-28T14:06:25-0500

The average product of labour is:

APL = q/L = 4KL^(- 0.5)/L = 4K/L^1.5 = 4×12×13^0.5 = 173 units.